Atomos Test Generator Free PYQs ExamsMy practice Log in / Sign up free

Probability (Class 12)

JEE Main 2026 (Online) 5th April Evening ShiftMCQEasy📈 Topic in 11% papers+4−1
Ques. 1JEE Main 2026, 5 April (Evening Shift)

The probabilities that players A and B of a team are selected for the captaincy for a tournament are 0.6 and 0.4, respectively. If A is selected the captain, the probability that the team wins the tournament is 0.8 and if B is selected the captain, the probability that the team wins the tournament is 0.7. Then the probability, that the team wins the tournament, is :

  1. 0.74
  2. 0.76
  3. 0.72
  4. 0.78

Answer

(2) 0.76

Solution

Let E1: A is the captain, E2: B is the captain, W: the team wins.

P(E1)=0.6, P(E2)=0.4, P(W E1)=0.8, P(W E2)=0.7

E1, E2 are mutually exclusive and exhaustive, by the theorem of total probability:

P(W)=P(E1)P(W E1)+P(E2)P(W E2)

P(W)=0.6 0.8+0.4 0.7=0.48+0.28=0.76

Answer: (2)

Keep practising — Probability (Class 12)

Ques. 2JEE Main 2026, 28 January (Morning Shift)

A bag contains 10 balls out of which k are red and (10 − k) are black, where 0 k 10. If three balls are drawn at random without replacement and all of them are found to be black, then the probability that the bag contains 1 red and 9 black balls is:

  1. 7110
  2. 711
  3. 755
  4. 1455
Ques. 3JEE Main 2026, 28 January (Evening Shift)

The probability distribution of a random variable X is given below : array|c|c|c|c|c|c|c|c|c| X & 4k & 307k & 327k & 347k & 367k & 387k & 407k & 6k \ P(X) & 215 & 115 & 215 & 15 & 115 & 215 & 15 & 115 \ array If E(X) = 26315, then P(X < 20) is equal to :

  1. 35
  2. 1415
  3. 815
  4. 1115
Ques. 4JEE Main 2026, 24 January (Morning Shift)

From a lot containing 10 defective and 90 non-defective bulbs, 8 bulbs are selected one by one with replacement. Then the probability of getting at least 7 defective bulbs is

  1. 73108
  2. 67108
  3. 7107
  4. 81108
Ques. 5JEE Main 2026, 23 January (Evening Shift)

Bag A contains 9 white and 8 black balls, while bag B contains 6 white and 4 black balls. One ball is randomly picked up from the bag B and mixed up with the balls in the bag A. Then a ball is randomly drawn from the bag A. If the probability, that the ball drawn is white, is pq, (p, q) = 1, then p + q is equal to

  1. 24
  2. 22
  3. 23
  4. 21
Ques. 6JEE Main 2026, 22 January (Morning Shift)

If a random variable x has the probability distribution arrayccccccccc x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 \ P(x) & 0 & 2k & k & 3k & 2k2 & 2k & k2 + k & 7k2 array then P(3 < x 6) is equal to

  1. 0.34
  2. 0.64
  3. 0.22
  4. 0.33
Ques. 7JEE Main 2026, 21 January (Evening Shift)

A random variable X takes values 0, 1, 2, 3 with probabilities 2a+130, 8a-130, 4a+130, b respectively, where a, b R. Let and respectively be the mean and standard deviation of X such that 2 + 2 = 2. Then ab is equal to:

  1. 12
  2. 60
  3. 30
  4. 3
Ques. 8JEE Main 2025, 8 April (Evening Shift)

If A and B are two events such that P(A) = 0.7, P(B) = 0.4 and P(A B) = 0.5, where B denotes the complement of B, then P (B (A B) ) is equal to

  1. 13
  2. 12
  3. 14
  4. 16
Ques. 9JEE Main 2025, 7 April (Evening Shift)

A bag contains 19 unbiased coins and one coin with head on both sides. One coin drawn at random is tossed and head turns up. If the probability that the drawn coin was unbiased, is mn, (m, n) = 1, then n2 - m2 is equal to :

  1. 64
  2. 80
  3. 60
  4. 72
Ques. 10JEE Main 2025, 7 April (Evening Shift)

Let a random variable X take values 0, 1, 2, 3 with P(X = 0) = P(X = 1) = p, P(X = 2) = P(X = 3) and E(X2) = 2E(X). Then the value of 8p - 1 is :

  1. 2
  2. 0
  3. 3
  4. 1
Ques. 11JEE Main 2025, 4 April (Morning Shift)

A box contains 10 pens of which 3 are defective. A sample of 2 pens is drawn at random and let X denote the number of defective pens. Then the variance of X is

  1. 1115
  2. 215
  3. 35
  4. 2875

All Probability (Class 12) PYQs →

PDFDownload Probability (Class 12) PYQs as PDF Free, with answer key and a QR code to every solution — print it or study offline⬇
All JEE Main Mathematics chapters